OpenStudy (anonymous):

If 3x2 – 2x + 7 = 0, then (x– 1 /3 ) 2 =

OpenStudy (nurali):

3x² - 2x + 7 = 0 3x² - 2x = - 7 x² - 2/3x = - 7/3 x² - 1/3x = - 7/3 + (1/3)² x² - 1/3x = - 21/9 + 1/9 (x - 1/3)² = - 20/9 Answer: (x - 1/3)² = - 20/9

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

My Pleasure.